OpenStudy (anonymous):

differentiate y = ln sqrt(e^3x (x^2 - 1))

OpenStudy (anonymous):

\[\ln \sqrt{e ^{3x}(x ^{2}-1)}\]

hartnn (hartnn):

use logarrithm properties. \(\ln \sqrt{x}=(1/2)ln x \\ ln(ab)=ln a+ ln b \\ ln e =1\)

OpenStudy (anonymous):

so first step would be \[\frac{ \ln (e ^{3x}) + \ln(x^{2}-1) }{ 2 }\]

hartnn (hartnn):

your first step is correct.

OpenStudy (anonymous):

what is the lne = 1? Do i then use the \[\ln f(x) = \frac{ f \prime (x) }{ f(x) }\] on the top two functions? What happens to the 2 at the bottom though?

hartnn (hartnn):

e is the base of natural logarithm, so logarithm of e = 1 ---> ln e =1 ln (e^(3x)) = 3x ln e = 3x. got this ?

OpenStudy (anonymous):

yes i see that now.

hartnn (hartnn):

keep bottom 2 as it is....till the end.

OpenStudy (anonymous):

so \[\frac{ 3x + \ln(x ^{2}-1) }{ 2 }\] is next step?

hartnn (hartnn):

yup. can u differentiate that? or need further help?

OpenStudy (anonymous):

ill try now if i cant ill post a msg.

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

is it \[\frac{ 3 + \frac{ 2x }{ x ^{2}-1 } }{ 2 }\]

hartnn (hartnn):

thats correct :) simplify it....

OpenStudy (anonymous):

\[\frac{ 3+ \frac{ 2x }{ (x+1)(x-1) } }{ 2 }\]

hartnn (hartnn):

i meant 3(x^2-1)+2x in the numerator.........

hartnn (hartnn):

\(\huge\frac{ 3 + \frac{ 2x }{ x ^{2}-1 } }{ 2 }=\frac{3(x^2-1)+2x}{2(x^2-1)}\) simplify the numerator.

OpenStudy (anonymous):

pardon?

OpenStudy (anonymous):

oh wait i see

OpenStudy (anonymous):

do I need to then simplify that further?

hartnn (hartnn):

yes, 3x^2+2x-3 in numerator.

hartnn (hartnn):

that would be it..... (3x^2+2x-3 ) / 2(x^2-1)

OpenStudy (anonymous):

thank you very much for your help. I got very confused as I have a program that suplies the answers and the answer to this question came to \[\frac{ \frac{ 3(x ^{2}-1)e ^{3x} }{ 2 } + xe ^{3x} }{ (x ^{2}-1)e ^{3x} }\]

OpenStudy (anonymous):

does that make any sense?

hartnn (hartnn):

i wonder how there still is e^(3x) term....

OpenStudy (anonymous):

thats mainly what confused me with this. Its Microsoft mathematics. Its worked for other sums but this one it went around the bend a little.

hartnn (hartnn):

is it explicitly mentioned to use product rule ??

OpenStudy (anonymous):

nope.

hartnn (hartnn):

i think they didn't simplify first, directly differentiated and used chain rule....

OpenStudy (anonymous):

is it wrong to do it that way?

hartnn (hartnn):

not wrong,but complicated....we found an equivalent answer in much simpler way.....

OpenStudy (anonymous):

yes thank you very much for that. Btw is it possible for you to explain the chain rule briefly?

hartnn (hartnn):

example : \(\large \frac{d}{dx}ln (sin x)=\frac{1}{sin x}\quad\frac{d}{dx}(sin x)=cos x/sinx=cot x.\)

OpenStudy (anonymous):

how does sin become cos/sin? surely f'(x) of sin = cos?

hartnn (hartnn):

sin did not become cos / sin derivative of ln(sin x) became cos / sin derivative of ln x = 1/ x so derivative of ln sin x = ?

OpenStudy (anonymous):

derivative of ln x = x'/x

hartnn (hartnn):

yup, thats chain rule.

OpenStudy (anonymous):

isnt chain rule f'(g(x)).g'(x)?

hartnn (hartnn):

same thing. g(x)=x,f(x)=ln x u just wrote the defination of derivative of f(g(x))

OpenStudy (anonymous):

ahhhhhh i see now. thank you. what if you have more than two functions though?

hartnn (hartnn):

f(g(h(x)))= f'(g(h(x))). g'(h(x)). h'(x)

OpenStudy (anonymous):

now that makes more sense than my lecturer lol.

hartnn (hartnn):

glad to hear that :)

OpenStudy (anonymous):

thank you very much.